# Personalizing Agent Privacy Decisions via Logical Entailment

### v2

@@ Abstract.
Personal language model-based agents are becoming more widespread for completing tasks on behalf of users; however, this raises serious privacy questions regarding whether these models will appropriately disclose user data. While prior work has evaluated language models on data-sharing scenarios based on general privacy norms, we focus on personalizing language models’ privacy decisions, grounding their judgments directly in prior user privacy decisions. Our findings suggest that general privacy norms are insufficient for effective personalization of privacy decisions. Furthermore, we find that eliciting privacy judgments from the model through In-context Learning (ICL) is unreliable to due misalignment with the user’s prior privacy judgments and opaque reasoning traces, which make it difficult for the user to interpret the reasoning behind the model’s decisions. To address these limitations, we propose ARIEL (Agentic Reasoning with Individualized Entailment Logic), a framework that jointly leverages a language model and rule-based logic for structured data-sharing reasoning. ARIEL is based on formulating personalization of data sharing as an entailment, whether a prior user judgment on a data-sharing request implies the same judgment for an incoming request. Our experimental evaluations on advanced models and publicly-available datasets demonstrate that ARIEL can reduce the F1 score error by 39.1% over language model-based reasoning (ICL), demonstrating that ARIEL is effective at correctly judging requests where the user would approve data sharing. Overall, our findings suggest that combining LLMs with strict logical entailment is a highly effective strategy for enabling personalized privacy judgments for agents.
@@ 1. Introduction
As Personal AI agents increasingly automate tasks, they frequently encounter scenarios requiring the transmission of private user data. Consequently, the agent must determine the appropriateness of such disclosures. Consider the restaurant booking scenario in Figure 1, where a restaurant requests the user’s phone number to finish booking a table. While the agent could consult the user, frequent interruptions lead to ’privacy fatigue,’ causing disengagement and ineffective data control from the user (acquisti2006there; choi2018role). Ideally, to alleviate this cognitive burden, the agent should be capable to autonomously judge when data sharing is appropriate.
Figure 1. High-level overview of a personal agent completing a task on behalf of a user. A third party is requesting user data in order to complete the task. The agent must decide whether it is appropriate to share the user’s information to fulfill the third-party’s request. To assist with the agent’s judgment, the agent is equipped with a knowledge base storing a history of prior user judgments on data-sharing requests.
@@
Prior to Large Language Models (LLMs), Personal Privacy Assistants (PPAs) relied on either rule-based (zhan2022model; zhan2023privacy) or traditional machine learning (ML) approaches (barbosa2019if; amoros2023predicting). These systems leverage a history of user prior user privacy judgments to decide whether to disclose data for incoming requests. For instance (Figure 1), if a user previously authorized sharing a phone number for a restaurant reservation, the agent utilizes this precedent to approve similar future requests. However, these methods face significant limitations: (1) rule-based approaches are domain-specific and difficult to transfer without substantial modification; and (2) ML methods require extensive user data (from many users) to achieve high performance.
@@
The integration of LLMs into agentic workflows (wang2024survey) offers a promising solution to these challenges, as their In-Context Learning (ICL) capabilities enable effective generalization even from sparse data (brown2020language). While recent studies have evaluated privacy judgments by LLMs (mireshghallah2024can; shao2024privacylens), they typically ground their definition of appropriateness in general Contextual Integrity (CI) norms derived from laws or societal consensus (nissenbaum2004privacy; nissenbaum2009privacy). However, relying solely on general norms fails to capture the nuances of user privacy decisions (barkhuus2012mismeasurement; guo2025not; tran2025understanding; zhang2025towards). In the example above, general norms would likely deem sharing a phone number for a restaurant reservation appropriate, yet a user with a restrictive history of data sharing might explicitly reject this request. General norms cannot account for individual privacy decisions, thus being a barrier in the development of personalized privacy agents. This critical gap motivates the central objective of our work.
@@
Objective: Personalizing LLM-based agents to make data-sharing decisions that maximize alignment with prior user decisions, minimize interruptions, and ensure auditability.
@@
  • Interpretability. The agent must provide an interpretable and traceable reasoning for every privacy judgment, allowing the user to audit the process and maintain trust in the agent’s decision-making.
  • User Agency. The agent must preserve user agency by identifying when prior user privacy judgments are insufficient or ambiguous, and escalating such requests to the user rather than forcing a judgment.
We observe that current LLM-based approaches fail to simultaneously satisfy these design requirements. While leveraging ICL on prior user privacy judgments improves personalization over general privacy norms, it does not guarantee alignment with prior user privacy decisions due to limited logical reasoning (lin2025zebralogic) and potential for hallucinations (ji2023survey; kalai2025language; zhang2023siren) from LLMs. Furthermore, the opaque nature of these models undermines interpretability by preventing verification of the reasoning steps used to reach a privacy judgment. Because the LLM-generated reasoning traces are often unfaithful to the actual decision process (barez2025chain; ye2022unreliability), users cannot effectively audit the agent, ultimately compromising their agency.
Thus, we propose ARIEL (Agentic Reasoning with Individualized Entailment Logic), a framework that satisfies the three requirements for personalization (Figure 3). The key insight is framing personalization of privacy judgments as an entailment, where the agent decides whether a user’s judgment on a prior request entails the same judgment for the incoming request. But rather than relying on LLMs to give a final appropriateness judgment, ARIEL combines LLMs reasoning and rule-based components for reliable and aligned reasoning. ARIEL leverages LLMs to generate ontologies that capture semantic relationships (e.g., mapping data sensitivity hierarchies), which are then processed by a rule-based component to provide a final judgment. This framework ensures that the agent acts only when a decision can be inferred by the user’s prior privacy judgments; otherwise, the request is escalated to the user.
We evaluate ARIEL against LLM-based reasoning with general privacy norms and prior user privacy judgments. These evaluations employ state-of-art LLMs – including GPT-5, Gemini 2.5 Pro, and Claude Sonnet 4.5 – on publicly available datasets covering smart home personal assistants (abdi2021privacy) and an educational (shvartzshnaider2016learning) sharing context. We find that (1) grounding privacy judgments in prior user privacy judgments yields significantly better personalization than relying solely on general privacy norms, and (2) ARIEL outperforms pure LLM-based reasoning by reducing the F1 score error of appropriate judgments by up to 39.1%. Our findings indicate that ARIEL is effective at correctly judging requests where the user would approve data sharing. We summarize our contributions below:
@@
    We demonstrate the limitations of aligning LLM agents using general privacy norms and ICL on prior user privacy judgments to make personalized data-sharing decisions on behalf of users.
@@
    We propose ARIEL, a framework for grounding the agent’s privacy judgments in prior user privacy judgments by logical entailment with LLM-generated ontologies. Jointly leveraging LLMs and logical entailment overcomes the rigidity of traditional rule-based systems without incurring the inconsistency inherent in pure LLM-based reasoning.
@@
    We demonstrate that ARIEL outperforms LLM-based reasoning for personalized privacy decisions, while ensuring user agency and interpretability. Notably, we find that performance gains are larger with smaller models, and that ARIEL remains effective with as few as 20 prior requests.
@@ 2. Related Work
We review prior work that relate to personalization of LLM-based agents to make privacy decisions aligned with user privacy preferences. We discuss (1) prior work in developing personal privacy assistants and (2) evaluations of privacy norms and personalization in LLMs. Finally, we examine works in (3) neurosymbolic reasoning and their relation to ARIEL.
@@ 2.1. Personal Privacy Assistants
Personal Privacy Assistants (PPAs) aim to reduce the user mental burden in making privacy decisions (stover2023investigating). They explore automating mobile app permissions, predicting user privacy preferences, and making information-sharing decisions (morel2025ai). Earlier PPAs targeting mobile app permissions clustered similar user privacy preferences to obtain privacy profiles that help determine mobile app permissions for a user (lin2014modeling; liu2016follow). Follow-up works developed a support vector machine model using a feature set of contextual cues, such as requesting app, the permission type, and the visibility of requesting app (wijesekera2018contextualizing). Other works use vignette surveys that capture contextual factors, such as device type, which serve as features for predicting privacy preferences using traditional ML techniques like logistic regression or collaborative filtering (naeini2017privacy; barbosa2019if; amoros2023predicting). Furthermore, researchers have proposed PPAs that manage data-sharing decisions by decomposing scenarios into Contextual Integrity (CI) information flows. These systems then utilize rule-based (zhan2022model; zhan2023privacy) or ML-based (kokciyan2022taking) approaches to infer the privacy decision. Finally, Natural Language Inference (NLI)-based models have been proposed to detect mismatches between different privacy policies or between policies and user privacy preferences (alshamsan2022detecting; hosseini2021ambiguity; rao2016expecting). While significant progress has been made for PPAs, they remain limited in that they require aggregating all user data for training, which requires users to upload their privacy preferences and prior decisions to a centralized entity. Moreover, these methods struggle with generalizing to different domains.
@@ 2.2. Contextual Privacy Awareness in LLMs
Recent work has evaluated LLMs on privacy-decision scenarios derived from general privacy norms, while also evaluating alignment of LLM privacy decisions in personalized settings.
Contextual Integrity (CI)
@@
Recently, researchers studied whether LLM can judge the appropriateness of data-sharing scenarios that are based on CI (nissenbaum2004privacy; nissenbaum2009privacy). In particular, mireshghallah2024can and shao2024privacylens developed ConfAIde and PrivacyLens, respectively, which are datasets that evaluate whether LLMs can reason about sharing data. yi2025privacy studied ambiguity in these datasets to incorporate contextual expansions for improved privacy reasoning. Furthermore, lan2025contextual improved LLM evaluation on PrivacyLens through post-training on a synthetic dataset. More recently, li2025privaci developed a benchmark targeting legal compliance. These datasets encompass privacy norms sourced from either legal statutes or aggregated judgments from crowdsource studies, both of which capture the general public’s privacy expectations on whether certain data sharing is appropriate in the particular context (guo2025not). Other works have operationalized the contextual awareness of LLMs to reduce privacy leakage of user tasks. ghalebikesabi2024operationalizing developed strategies for LLM-based assistants to be CI compliant in form filling tasks. bagdasarian2024airgapagent implemented a minimization module with an LLM to generate a minimized user data set based on the user task. ngong2025protecting utilized LLMs to reduce private information within a user’s prompt. While evaluating LLMs on general privacy norms is important, these norms fail to capture the nuances of individual user preferences (barkhuus2012mismeasurement; guo2025not; tran2025understanding; zhang2025towards). Consequently, reliance on general norms alone is ineffective for developing agents tasked with making personalized privacy decisions.
Personalized Settings
@@
groschupp2025can investigated personalized access control in LLMs, where the LLM must determine whether to grant data access by evaluating an incoming request against the user’s stated access-control preferences (e.g. ”I prefer to not share data unless necessary”). However, it might not be practical to assume that users can explicitly specify access-control preferences. Further, delegating the access control decision to LLMs can struggle to align with the user’s access-control preferences, especially with limitations of LLMs such hallucinations (ji2023survey; kalai2025language; zhang2023siren) and biases (gallegos2024bias). Another work studied how well LLMs align with human preferences in household human-robot interaction contexts (sullivan2025benchmarking). Our work, on the other hand, focuses on a different problem setup, where we evaluate LLMs making privacy decisions on behalf of users. The final work evaluates how well do judgments from LLMs align with user expectations (cheng2024ci). However, their evaluation is quite limited as the dataset is synthetically generated with the author’s annotating the user appropriateness labels.
@@ 2.3. Neurosymbolic Approaches for Reasoning
Prior works that are methodologically similar to ours broadly fall under neurosymbolic reasoning, specifically where the LLM acts as an interface to user inputs and offloads the reasoning through rule-based/logical approaches. Such works include using an LLM to translate natural language into a formal intermediate language, which is then offloaded to a deterministic solver to perform the logical reasoning (olausson2023linc; pan2023logic; trinh2024solving; ye2023satlm). While ARIEL does not employ a theorem solver, it does use few-shot prompting with LLMs to map data-sharing requests to levels in a corresponding ontology, which can then use a rule-base component to determine entailment.
@@ 3. Personalizing Privacy Decisions
As illustrated in Figure 1, our agent setup involves a personal agent performing tasks on a user’s behalf. The agent has access to the user’s data and manages data-sharing requests from third parties required to complete these tasks. We describe below our problem setup of how the agent decides whether it is appropriate to transmit user data.
@@ 3.1. Problem Statement
We address the problem of an LLM-based agent $A_{\theta}$ tasked with making personalized privacy decisions for a user $u$. We assume a non-adversarial setting where the requests $R_{u}$ are not maliciously designed to manipulate the agent $A_{\theta}$. Upon receiving a request $R_{u}$ to share data $d_{u}$, the agent must infer the user’s judgment $l_{u}$ by leveraging a knowledge base $D_{u}$:
$A_{\theta}(D_{u},R_{u}\in\{\text{appropriate, inappropriate}\}.$
The knowledge base $D_{u}$ represents a history of the user’s past privacy judgments $l^{i}_{u}$ on prior requests $R^{i}_{u}$:
@@
Let the incoming request $R_{u}$ be a bank asking for the user’s partial Social Security Number (SSN) to open a checking account. The knowledge base $D_{u}$ contains a prior request $R^{i}_{u}$ where the same bank asked for the user’s full SSN for the same purpose, which the user judged as appropriate ($l^{i}_{u}=\texttt{appropriate}$). The agent must decide if it can infer the appropriateness of $R_{u}$ based on this precedent $(R^{i}_{u},l^{i}_{u})$.
@@
Using Definition 3.2, we can map the prior and incoming request from Example 3.1 to the corresponding five parameters, as the following:

### v3

@@ Abstract.
Personal large language model (LLM) agents increasingly perform tasks that require access to user data, raising concerns about appropriate data disclosure. We show that relying solely on LLMs to make data-sharing decisions is insufficient. Prompting LLMs with general privacy norms fails to capture individual users’ privacy preferences, while providing prior user data-sharing decisions through in-context learning (ICL) leads to unreliable and opaque reasoning. To address these limitations, we propose ARIEL (Agentic Reasoning with Individualized Entailment Logic), a framework that combines LLMs with rule-based logic to enable structured, personalized privacy reasoning. The core mechanism of ARIEL determines whether a user’s prior decision on a data-sharing request logically entails the same decision for a new request. Experimental evaluations using advanced models and public datasets show that ARIEL reduces the F1 error rate for appropriate judgments by 40.6% compared to standard ICL-based reasoning, indicating that ARIEL is effective at correctly judging requests where the user would approve data sharing. These results demonstrate that integrating LLMs with logical entailment provides an effective and interpretable approach for automating personalized privacy decisions.
@@ 1. Introduction
As personal AI agents increasingly automate tasks, they frequently encounter scenarios requiring the transmission of private user data. Consequently, the agent must determine the appropriateness of such disclosures. Consider the restaurant booking scenario in Figure 1, where a restaurant requests the user’s phone number to complete the table reservation. While the agent could seek the user’s permission to transmit their phone number, frequent interruptions lead to ’privacy fatigue,’ causing disengagement and ineffective data control by the user (acquisti2006there; choi2018role). To alleviate this cognitive burden, the agent should ideally be capable of autonomously judging when data sharing is appropriate.
@@
Prior to Large Language Models (LLMs), Personal Privacy Assistants (PPAs) relied on either rule-based (zhan2022model; zhan2023privacy) or traditional machine learning (ML) approaches (barbosa2019if; amoros2023predicting). These systems leverage a history of prior user privacy judgments to decide whether to disclose data for incoming requests. For instance, as shown in Figure 1, if a user previously authorized sharing a phone number for a restaurant reservation, the agent can utilize this precedent to approve similar future requests. However, these methods face significant limitations: (1) rule-based approaches are domain-specific and difficult to transfer without substantial modification; and (2) ML methods require extensive user data (from many users) to achieve high performance.
Figure 1. A schematic visualization of the problem setup with ARIEL, our proposed privacy-reasoning framework, operating in a larger end-to-end agentic system. ① A user sends a task to the agent to complete. ② A third party requests user data in order to complete the task. ③ After the data request has been authenticated, the agent retrieves the relevant user data and prior user judgments on data-sharing requests from memory. ④ The agent decides whether it is appropriate to share the user’s information to fulfill the third party’s request. ARIEL either allows or denies transmitting the user data, or escalates the data-sharing request to the user for further assistance.
@@
The integration of LLMs into agentic workflows (wang2024survey) offers a promising solution to these challenges, as their In-Context Learning (ICL) capabilities enable effective generalization even from sparse data (brown2020language). While recent studies have evaluated privacy judgments by LLMs (mireshghallah2024can; shao2024privacylens), they typically ground their definition of appropriateness in general Contextual Integrity (CI) norms derived from laws or societal consensus (nissenbaum2004privacy; nissenbaum2009privacy). However, relying solely on general norms fails to capture the nuances of user privacy decisions (barkhuus2012mismeasurement; guo2025not; tran2025understanding; zhang2025towards). In the example above, it is generally acceptable to share a phone number for a restaurant reservation, yet a user with a restrictive history of data sharing might explicitly reject this request. General norms cannot account for individual privacy decisions, thus being a barrier in the development of personalized privacy agents. This critical gap motivates the central objective of our work.
@@
Objective: Personalizing LLM-based agents to make data-sharing decisions that maximize alignment with prior user decisions, minimize user intervention, and ensure auditability.
@@
  • Interpretability. The agent must provide interpretable and traceable reasoning for every privacy judgment, allowing the user to audit the process and maintain trust in the agent’s decision-making.
  • User Agency. The agent must preserve user agency by identifying when prior user privacy judgments are insufficient or ambiguous then escalating such requests to the user rather than forcing a judgment.
We observe that current LLM-based approaches fail to simultaneously satisfy these design requirements. While leveraging ICL on prior user privacy judgments improves personalization over general privacy norms, it does not achieve effective alignment with prior user privacy decisions due to limited logical reasoning (lin2025zebralogic) and potential for LLM hallucinations (ji2023survey; kalai2025language; zhang2023siren). Furthermore, the opaque nature of these models undermines interpretability by preventing verification of the reasoning steps used to reach a privacy judgment. Because LLM-generated reasoning traces are often unfaithful to the actual decision process (barez2025chain; ye2022unreliability), users cannot effectively audit the agent, ultimately compromising their agency.
Thus, we propose ARIEL (Agentic Reasoning with Individualized Entailment Logic), a framework that satisfies these three requirements. Our key insight is to frame personalized privacy judgments as logical entailment: determining whether a user’s judgment on a prior request entails the same judgment for an incoming request. Rather than relying on LLMs to provide an opaque appropriateness judgment, ARIEL combines LLM-based reasoning and rule-based components for reliable and aligned judgments. Specifically, ARIEL leverages an LLM to generate ontologies that capture semantic relationships from the user’s prior privacy judgments (e.g., a data sensitivity hierarchy), which are then processed by a rule-based component to provide a final judgment. This framework ensures that the agent acts only when a decision can be inferred from the user’s prior privacy judgments; otherwise, the request is escalated to the user (Figure 1).
We evaluate ARIEL against LLM-based reasoning baselines that rely on either general privacy norms or prior user privacy judgments. These evaluations were conducted using state-of-the-art LLMs—including GPT-5, Gemini 2.5 Pro, and Claude Sonnet 4.5—on publicly available datasets covering smart home assistant (abdi2021privacy) and educational (shvartzshnaider2016learning) data-sharing contexts. We find that (1) grounding privacy judgments in prior user privacy judgments yields significantly better personalization than relying solely on general privacy norms, and (2) ARIEL outperforms pure LLM-based reasoning by reducing the F1 score error of appropriate judgments by up to 40.6%. Our findings indicate that ARIEL is effective at correctly judging requests where the user would approve data sharing. We summarize our contributions below:
@@
    We demonstrate the limitations of aligning LLM agents with general privacy norms, as well as the shortcomings of current LLM-based approaches that use prior user judgments for personalized data-sharing decisions.
@@
    We propose ARIEL, a framework for grounding the agent’s privacy judgments in prior user privacy judgments through logical entailment with LLM-generated ontologies. Jointly leveraging LLMs and logical entailment overcomes the rigidity of traditional rule-based systems without incurring the inconsistency inherent in pure LLM-based reasoning.
@@
    We show that ARIEL outperforms LLM-based reasoning for personalized privacy decisions, while incorporating user agency and interpretability. Notably, we find that performance gains are larger with smaller models (Gemma 3 4B) and that ARIEL remains effective with as few as 20 prior requests.
@@ 2. Related Work
We review closely-related work on personalizing LLM-based agents to make privacy decisions aligned with user privacy preferences. We discuss (1) prior work in developing personal privacy assistants and (2) evaluations of privacy norms and user preferences in LLMs. Finally, we examine works in (3) neurosymbolic reasoning and their relation to ARIEL.
@@ 2.1. Personal Privacy Assistants
Personal Privacy Assistants (PPAs) aim to reduce the user mental burden in making privacy decisions (colnago2020informing; stover2023investigating). They explore automating mobile app permissions, predicting user privacy preferences, and making information-sharing decisions (morel2025ai). Earlier PPAs targeting mobile app permissions clustered similar user privacy preferences to derive privacy profiles that help determine mobile app permissions for individual users (lin2014modeling; liu2016follow). Follow-up works utilized contextual factors, such as device type, requesting app, and the permission type, as features for predicting privacy preferences using traditional ML techniques like logistic regression or collaborative filtering (naeini2017privacy; barbosa2019if; amoros2023predicting; wijesekera2017feasibility; wijesekera2018contextualizing). Other works have proposed PPAs that manage data-sharing decisions by decomposing scenarios into Contextual Integrity (CI) information flows. These systems then utilize rule-based (zhan2022model; zhan2023privacy) or ML-based (kokciyan2022taking) approaches to infer the privacy decision. Finally, Natural Language Inference (NLI)-based models have been proposed to detect mismatches between different privacy policies or between policies and user privacy preferences (alshamsan2022detecting; hosseini2021ambiguity; rao2016expecting). While significant progress has been made for PPAs, they remain limited in that they require aggregating all user data for training, which requires users to upload their privacy preferences and prior decisions to a centralized entity. Moreover, these methods struggle with generalizing to different domains.
@@ 2.2. Contextual Privacy Awareness in LLMs
Recent work has evaluated LLMs on privacy-decision scenarios derived from general privacy norms, while also evaluating alignment of LLM privacy decisions with user privacy preferences.
General Privacy Norms
@@
Recently, researchers studied whether LLMs can judge the appropriateness of data-sharing scenarios that are based on CI (nissenbaum2004privacy; nissenbaum2009privacy). In particular, mireshghallah2024can and shao2024privacylens developed ConfAIde and PrivacyLens, respectively, which are datasets that evaluate whether LLMs can reason about sharing data. yi2025privacy studied ambiguity in these datasets to incorporate contextual expansions for improved privacy reasoning. Furthermore, lan2025contextual improved LLM evaluation on PrivacyLens through post-training on a synthetic dataset. More recently, li2025privaci developed a benchmark targeting legal compliance. And mireshghallah2025cimemories evaluates CI compliance of LLMs with persistent memory that was synthetically generated. These datasets encompass privacy norms sourced from either legal statutes or aggregated judgments from crowdsourced studies, both of which capture the general public’s privacy expectations on whether certain data sharing is appropriate in a particular context. Other works have operationalized the contextual awareness of LLMs to reduce privacy leakage of user tasks. ghalebikesabi2024operationalizing developed strategies for LLM-based assistants to be CI compliant in form filling tasks. bagdasarian2024airgapagent implemented a minimization module with an LLM to generate a minimized set of data based on the user’s task. ngong2025protecting utilized LLMs to reduce private information within a user’s prompt. While evaluating LLMs on general privacy norms is important, these norms fail to capture the nuances of individual user preferences (barkhuus2012mismeasurement; guo2025not; tran2025understanding; zhang2025towards). Consequently, reliance on general norms alone is ineffective for developing agents tasked with making personalized privacy decisions.
User Privacy Preferences
@@
groschupp2025can and wu2025towards investigated personalized access control in LLMs, where the LLM must determine whether to grant data access by evaluating an incoming request against a user’s access-control preferences (e.g., ”I prefer not to share data unless necessary”). However, it might not be practical to assume that users can explicitly specify access-control preferences. Further, delegating the access control decision to LLMs may result in misalignment with the user’s access-control preferences, especially with limitations of LLMs such as hallucinations (ji2023survey; kalai2025language; zhang2023siren) and biases (gallegos2024bias). Another work studied how well LLMs align with human preferences in household human-robot interaction contexts (sullivan2025benchmarking). Our work, on the other hand, focuses on a different problem setup, where we evaluate LLMs making privacy decisions on behalf of users. The final work evaluates how well judgments from LLMs align with user expectations (cheng2024ci). However, their evaluation is quite limited as the dataset is synthetically generated with the authors annotating the user-appropriateness labels.
@@ 2.3. Neurosymbolic Approaches for Reasoning
Prior works that are methodologically similar to ours broadly fall under neurosymbolic reasoning, specifically where the LLM acts as an interface to user inputs and offloads the reasoning to rule-based/logical approaches. Such works include using an LLM to translate natural language into a formal intermediate language, which is then offloaded to a deterministic solver to perform the logical reasoning (olausson2023linc; pan2023logic; trinh2024solving; ye2023satlm). While ARIEL does not employ a theorem solver, it does use few-shot prompting with an LLM to map data-sharing requests to levels in a corresponding ontology, which are then passed through a rule-base component to determine entailment.
@@ 3. Personalizing Privacy Decisions
As illustrated in Figure 1, a personal agent executes tasks on a user’s behalf. The agent has access to the user’s data and manages data-sharing requests from third parties required to complete these tasks. Below, we formalize the problem setup describing how the agent determines whether transmitting user data is appropriate.
@@ 3.1. Problem Statement
We consider an LLM-based agent $A_{\theta}$ that leverages persistent memory $D_{u}$ (chhikara2025mem0; xu2025mem) to make personalized privacy decisions for a user $u$. We assume a non-adversarial setting where the data-sharing requests $R_{u}$ are either authenticated or otherwise not maliciously designed to manipulate the agent $A_{\theta}$. Upon receiving a request $R_{u}$ to share data $d_{u}$, the agent must infer the user’s judgment $l_{u}$:
$A_{\theta}(D_{u},R_{u})\in\{\text{appropriate, inappropriate}\}.$
The memory $D_{u}$ represents a history of the user’s past privacy judgments $l^{i}_{u}$ on data-sharing requests $R^{i}_{u}$:
@@
Let the incoming request $R_{u}$ be from a bank asking for the user’s partial Social Security Number (SSN) to open a checking account. The memory $D_{u}$ contains a prior request $R^{i}_{u}$ in which the same bank asked for the user’s full SSN for the same purpose, which the user judged as appropriate ($l^{i}_{u}=\texttt{appropriate}$). The agent must decide whether it can infer the appropriateness of $R_{u}$ based on this precedent $(R^{i}_{u},l^{i}_{u})$.
@@
Using Definition 3.2, we can map the prior and incoming requests from Example 3.1 to the corresponding five parameters, shown in Figure 2.
@@
Figure 2. A example mapping of the prior request with user judgment and the incoming request from Example 3.1 to the five parameters.